OpenStudy (anonymous):

Medal A hair dye is made by blending a 7% hydrogen peroxide solution and a 4% hydrogen peroxide solution. How many milliliters of each are used to make a 300- milliliters solution that is 5% hydrogen peroxide?

OpenStudy (anonymous):

x will be the amount of 7% solution so, 0.07x+0.04(300-x)=0.05(300) 0.07x+12-0.04x=15 0.07x+12-12-0.04x=15-12 0.03x=3 x=100 ml of 7% solution is needed 300-x = 300-100 = 200 ml of 4% solution is needed

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

yes i understand it now thank you :)